atoms of Carbon dioxide gas decay from upper level to lower level with a spread in frequency is about 10kHz. calculate the line shape if the broadening mechanism is found Doppler broadening?
1
Expert's answer
2020-05-21T13:35:22-0400

The Gaussian lineshape is


"G(\omega)=\frac{1}{\Gamma\sqrt{\pi}}e^{-\frac{(\omega-\omega_0)^2}{\Gamma^2}}."